# Louis Pallegoix1, Jaime Travesedo1, Alexandre S. May1,2, Léo Balembois1, Denis Vion1, Patrice Bertet*1 and Emmanuel Flurin Université Paris-Saclay, CEA, CNRS, SPEC, 91191 Gif-sur-Yvette Cedex, France 2Alice&\&&Bob, 53 boulevard du Général Martial Valin, 75015 Paris

### v1

@@ IV.2 Four-Wave-Mixing tuning
The Four-Wave mixing condition is found by sending a pump tone at frequency $\omega_{p}$ together with a signal at frequency $\omega$ into the SMPD input, followed by qubit readout. Sweeping $\omega_{p}$ and $\omega$, the 4WM condition is identified by an enhanced qubit excited state when the signal is present, which vanishes when it is switched off (see Fig. 4). The pump power is then tuned to reach unit cooperativity (see Appendix and Fig. 4).
Figure 4: a) Four-wave mixing pattern : this color plot shows the measured qubit excited state as a function of pump frequency $\omega_{p}$ and signal frequency $\omega$. b) Fit from a two-coupled cavity model. To fit the data, the value of the cooperativity $C$ is adjusted. Here, the value is $C=0.99$ (see Appendix A). Other fit parameters are the buffer linewidth, $\nicefrac{{\kappa_{b}}}{{2\pi}}=120$ kHz and waste linewidth, $\nicefrac{{\kappa_{w}}}{{2\pi}}=1.75$ MHz. c) Same experiment as a with the signal tone switched off. The qubit is excited only when the signal is close to $\omega_{b}$, and the pump verifies the $4$-Wave Mixing condition. Insets : corresponding pulse sequences. Purple pulses are sent through the pump/drive line, orange pulses through the signal/input line and green pulses through the waste/readout (RO) line. Pumping, signal and readout pulses are respectively 15 µs, 10.5 µs and 800 ns long.
@@ IV.3 Cyclic operation : Detector Bandwidth and Dark counts
We now turn to the SMPD operation in cyclic mode. The detection step duration is set to $T_{d}=15$ µs. When no photon is sent to the device, the qubit is most often in the ground state at the end of the sequence, implying that no reset is needed, and that the average cycle duration is very close to $T_{d}+T_{\mathrm{RO}}=15.8$ µs, yielding the duty cycle $\eta_{\mathrm{cycle}}=0.95$. From Eq. 3, we expect $\eta_{\mathrm{SMPD}}$ to reach $0.74$.
@@
The experimental $\eta_{\mathrm{SMPD}}$ is measured by counting the number of clicks detected over 1 second integration time, at the same time that a calibrated signal is sent to the SMPD input. The signal power is calibrated using the qubit as a detector (see [1] and Appendix B). The result is plotted as a function of input frequency $\omega$ in Fig. 2d. The maximum efficiency reaches $\eta_{\mathrm{SMPD}}=0.8$. This is even slightly higher than the predicted value, which implies in particular that the Four-Wave-Mixing efficiency $\eta_{\mathrm{4WM}}$ is very close to $1$. Fitting $\eta_{\mathrm{SMPD}}(\omega)$ by a Lorentzian, we obtain the detector bandwidth $\kappa_{d}$. Two curves are shown, obtained for two different values of $\kappa_{b}$. The maximum efficiency is unchanged, but the detector bandwidth is changed according to $\kappa_{b}$. This confirms that the SMPD bandwidth can be adjusted by tuning $\kappa_{b}$ via the Purcell filter frequency. We measure $\kappa_{d}/2\pi=315$ kHz for $\kappa_{b}/2\pi=170$ kHz, and $\kappa_{d}/2\pi=530$ kHz for $\kappa_{b}/2\pi=280$ kHz, thus confirming the approximate relation $\kappa_{d}\simeq 2\kappa_{b}$ (see Appendix A).
@@
Dark counts are measured by recording photon counting traces in the absence of any input signal (see Table 3, where the detector bandwidth was $\kappa_{d}/2\pi=170$ kHz). To obtain the various contributions to the dark counts, the traces are systematically measured with pump off (yielding $\alpha_{q}$), pump on detuned by $20$ MHz (yielding $\alpha_{q}+\alpha_{p}=\alpha_{\mathrm{err}}$), and pump on tuned to the 4WM conditions (yielding $\alpha=\alpha_{\mathrm{err}}+\alpha_{\mathrm{th}}$). At $10$ mK, we obtain $\alpha_{q}=8\mathrm{s}^{-1}$, $\alpha_{p}=2\mathrm{s}^{-1}$, and $\alpha_{\mathrm{th}}=21\mathrm{s}^{-1}$. Combined with the measured efficiency, we obtain a power sensitivity $S=3\cdot 10^{-23}\mathrm{W}/\sqrt{\mathrm{Hz}}$, a factor of three better than the previous state-of-the-art [1], which corresponds to an order of magnitude gain in measurement time.
@@
Time traces are similarly measured by varying the cryostat temperature $T$. The values of $\alpha_{q}$ and $\alpha_{\mathrm{th}}$ are shown as a function of $T$ in Fig. 5. The qubit contribution increases with temperature, because of the higher excited state probability at equilibrium; it is well fitted by $\alpha_{q,0}+K_{q}p_{\mathrm{th},q}(T)$, with $\alpha_{q,0}=7\,\mathrm{s}^{-1}$ and $K_{q}=2.2\cdot 10^{4}\,\mathrm{s}^{-1}$. The value of $K_{q}$ is slightly larger than the predicted $(T_{d}/T_{1})(1/(T_{d}+T_{\mathrm{RO}}+T_{\mathrm{reset}}))$ (see Section IIIB), which is $1.4\cdot 10^{4}\,\mathrm{s}^{-1}$, possibly because of a reduction of $T_{1}$ at higher temperature. The thermal contribution also increases with temperature; it is well fitted by $\alpha_{\mathrm{th},0}+K_{\mathrm{th}}\bar{n}_{\mathrm{th},b}$, with $\alpha_{\mathrm{th},0}=31\mathrm{s}^{-1}$, and $K_{\mathrm{th}}=2\cdot 10^{5}\,\mathrm{s}^{-1}$. This value is in quantitative agreement with the expected $\eta_{\mathrm{SMPD}}\kappa_{d}/4=2.1\cdot 10^{5}\mathrm{s}^{-1}$, thus providing an independent confirmation of the efficiency estimate reported earlier.
@@
The equilibrium qubit excitation (without active reset) $p_{\mathrm{th},q}$ at $10$ mK translates into an effective qubit temperature of $42$ mK. The thermal photon rate $\alpha_{\mathrm{th},0}$ measured at $10$ mK translates into a mean photon number per mode of $1.5\cdot 10^{-4}$ and thus to an effective temperature of $40$ mK. This is higher than the mean photon number expected from our line attenuation, which we estimate to be below $\sim 2\cdot 10^{-5}$. Our results therefore confirm the well-known difficulty to thermalize microwave fields and modes to temperatures much lower than $\sim 40$ mK, possibly due to the difficulty of perfectly thermalizing microwave components. It is interesting to note that SMPD devices are ideally well suited to study this phenomenon, which represents a major issue for superconducting quantum computers.
@@
Table 3: Photon counting traces taken in different conditions. First line : Mixing chamber temperature is 10 mK and the pump is off, yielding $\alpha_{q}$. Second line : Mixing chamber temperature is 10 mK and pump is switched on, but $20$ MHz detuned from the four-wave mixing condition. This yields $\alpha_{\mathrm{err}}=\alpha_{q}+\alpha_{p}$. Third line : Mixing chamber temperature is 10 mK and the pump is on at the 4WM condition, yielding $\alpha_{\mathrm{err}}+\alpha_{\mathrm{th}}$. Three following lines : Mixing chamber temperature is risen and the count rate rises accordingly, as detailed on fig. 5.
@@ IV.4 Dependence on detection bandwidth
Figure 6: Blue full circles are the measured total dark counts, $\alpha_{\mathrm{err}}+\alpha_{\mathrm{th}}$, as a function of SMPD bandwidth $\kappa_{d}$. Dashed line is a linear fit, with slope $1.5\cdot 10^{-5}$, and intercept $11.6\mathrm{s}^{-1}$. Red triangles show the measured efficiency at resonance, $\eta_{\mathrm{SMPD}}$.
We finally study the SMPD figures of merit as a function of detection bandwidth, at the lowest temperature of $10$ mK. We measure the maximum value of $\eta_{\mathrm{SMPD}}(\omega)$ and $\alpha$ as a function of $\kappa_{d}$ (see Fig. 6). We observe that $\eta_{\mathrm{SMPD}}$ increases with $\kappa_{d}$, although the dependence is very weak as soon as the SMPD bandwidth is above $250$ kHz. This is expected since the measurements are performed with a monochromatic tone.The reduction in efficiency at the lowest value of $\kappa_{d}$ may be attributed to the contribution of internal buffer losses, as well as possible transient effects that effectively shorten the detection time $T_{d}$.
The false positive rate $\alpha$ on the other hand scales linearly with $\kappa_{d}$, as expected from Eq. 5. The fitted slope is $1.6\cdot 10^{-5}$, in agreement with the expected $\eta\bar{n}_{\mathrm{th},b}/4$ yielding a thermal mode occupation $\bar{n}_{\mathrm{th},b}=0.8\cdot 10^{-4}$, slightly lower than in the measurements of Fig. 5 and corresponding to a mode temperature of $38$ mK. The intercept at $\kappa_{d}=0$ is $11.6\,\mathrm{s}^{-1}$, consistent with the error rate $\alpha_{\mathrm{err}}$ measured with pump detuned. Overall, the oberved linear dependence of $\alpha_{\mathrm{th}}$ on $\kappa_{d}$ proves that these counts are indeed due to thermal photons impinging on the SMPD input line. Not much can be inferred about their spectrum, apart from the fact that it appears to be flat on the scale of the detection bandwidth ($\sim 1$ MHz).
@@ IV.5 Single-spin measurements
We confirm the high efficiency of the detector by measuring the microwave fluorescence from a single electronic spin. The setup closely follows refs  [8] and  [38], and is schematically depicted in Fig. 7. The electronic spin is an $\mathrm{Er}^{3+}$ ion in a $\mathrm{CaWO}_{4}$ matrix and is coupled to a superconducting resonator to enhance its radiative decay rate $\Gamma_{R}$. The spin is excited by a $\pi$ pulse with near unit probability. It then relaxes radiatively by emitting a photon, which is either absorbed by the spin resonator internal losses, or emitted in the line, after which it may reach the SMPD detector or be absorbed by the losses in-between the two devices. The SMPD buffer $\omega_{b}$ is tuned to the spin and resonator frequency. The spin-to-click probability is therefore $\eta=\eta_{reso}\eta_{loss}\eta_{\mathrm{SMPD}}(\omega_{b})$, with $\eta_{reso}=0.6$ in this device (described in more detail elsewhere [38]).
The spin microwave fluorescence curve is shown in Fig. 7. It is well fitted by an exponential decay with time constant $\Gamma_{R}^{-1}=1.24$ ms on top of a constant background, due to the SMPD dark counts. From the integral below the fluorescence curve, we obtain $\eta=0.4$. This implies that the microwave power absorption in-between the spin device and the SMPD is $\eta_{loss}=0.85$, a very plausible value given that the line includes a superconducting cable, a circulator, and an infra-red filter. This provides an independent confirmation of the high efficiency reached by this SMPD device and of its interest for single-spin spectroscopy.
@@ V Conclusion
We have demonstrated a microwave photon counter with tunable frequency (in the $7$ GHz range) and bandwidth (between $\sim 100$ kHz and $\sim 1$ MHz). The device has a maximum efficiency of $0.8$, and its noise can be reduced by lowering the bandwidth. At an input bandwidth of $200$ kHz, we demonstrate a record power sensitivity of $3\cdot 10^{-23}\mathrm{W}/\sqrt{\mathrm{Hz}}$. The device is well suited to detect emitters with a narrow line, and will find applications in single-spin magnetic resonance [38, 39] as well as haloscope-based dark matter search [36].
@@ Acknowledgements
We acknowledge technical support from P. Sénat, D. Duet, P.-F. Orfila and S. Delprat, and are grateful for fruitful discussions within the Quantronics group. We acknowledge support from the Agence Nationale de la Recherche (ANR) through the MIRESPIN (ANR-19-CE47-0011) project. We acknowledge support of the Région Ile-de-France through the DIM QUANTIP, from the AIDAS virtual joint laboratory, and from the France 2030 plan under the ANR-22-PETQ-0003 grant. This project has received funding from the European Union Horizon 2020 research and innovation program under the project OpenSuperQ100+, and from the European Research Council under the grant no. 101042315 (INGENIOUS). We acknowledge IARPA and Lincoln Labs for providing the Josephson Traveling-Wave Parametric Amplifier.
Figure 7: Microwave fluorescence of an individual $\mathrm{Er}^{3+}$ electron spin in a $\mathrm{CaWO}_{4}$ host crystal matrix at $10$ mK, detected by the SMPD reported here operated with $\kappa_{d}/2\pi=500$ kHz. Top : scheme of the experiment. The $Er^{3+}$ spin is magnetically coupled to the microwave field $\vec{B}_{1}$ of an LC resonator and Zeeman tuned to resonance by a field $||\vec{B}_{0}||$ = 460 mT. The spin is excited by a $\pi$ pulse, and the photon it emits upon radiative relaxation is detected by the SMPD. Graph : blue full circles are the measured count rate following a $\pi$ pulse excitation of the spin, as a function of the time. Solid blue line is an exponential fit, from which we extract the spin lifetime $\Gamma_{R}^{-1}=1.24$ ms and spin-to-click efficiency, $\eta=0.4$. Note that the noise floor is slightly higher than expected from Fig. 6, possibly due to a different wiring scheme in the cryostat.
@@
In this section we detail the pump parameters tuning procedure. The optimal pump frequency should satisfy the 4WM condition, $\omega_{\mathrm{4WM}}=\omega_{q}+\omega_{w}-\chi_{w}-\omega_{b}$.
The pump amplitude is chosen by recording two-dimensional plots such as shown in Fig. 4 of the main text, where the qubit excited state probability is recorded as a function of pump and signal frequency, $\omega_{p}$ and $\omega$, respectively, for various pump powers. Each plot is then fitted to the following formula [40] :
@@
with $\delta=\omega-\omega_{b}$ is the signal-buffer detuning, and $\delta_{p}=\omega_{p}-\omega_{\mathrm{4WM}}$ the pump detuning from the 4WM condition.
@@
Measuring the SMPD efficiency starts by estimating the power at the buffer input. This is achieved by using the qubit as a detector. Ramsey fringes are shifted by $\Delta_{q}$ and dephased by $\Gamma_{q}$ according to the equation:
@@
with $\epsilon$ the complex amplitude of the intra-cavity field. The efficiency is then simply computed by dividing the number of clicks per second by the number of photons per second at the device input.
@@
The process starts from a two-inch, double-side polished sapphire wafer. This wafer is dipped in a piranha solution made of 2/3 of sulfuric acid and 1/3 of hydrogen peroxyde (in volume) during 20 minutes. Then, it is rinsed with water and dried by dinitrogen blow. After that, the wafer is loaded into an RF reactive sputterer together with a tantalum target. The targeted thickness of tantalum sputtering is between 80 and 90 nm. The sputtering process shall be performed while the wafer is heated up to 600°C in order to get alpha phase tantalum.
The wafer is then prepared for dicing. In that regard, UV III is spincoated over the wafer during 60 seconds at 3000 rpm with an acceleration of 2000 rpm/s. This is followed by a postbake at 120°C during 3 minutes.
@@
After dicing, the protective UV III layer is removed by shooting IPA with a pissette on the chosen chip during 15 seconds, then again during 5 seconds. For further cleaning, the chip is dipped during 5 minutes in a bath of acetone with strong ultrasounds, then in another bath of isopropanol (IPA) during 3 minutes, still with strong ultrasounds. In the following, it is dried by dinitrogen blow and dipped in a piranha solution during 20 minutes. The chip is then rinsed with water, dehydrated with IPA, dried by dinitrogen blow and further dehydrated on a hotplate at 120°C during 5 minutes. Then comes the spincoating of AZ1518 during 60 seconds at 3000 rpm with an acceleration of 2000 rpm/s. The postbake is performed on hotplate during 2 minutes at 115°C. The circuit is then patterned using a UV LASER direct writer. A supplementary exposure is needed at the edge bead. The development is performed in pure MF319 during 60 seconds and stopped with water. The chip is dried by dinitrogen blow and hardbaked on hotplate during 2 minutes at 115°C. Eventually, the circuit is etched by dipping in Transene 111 (1/3 nitric acid, 1/3 hydrofluoric acid and 1/3 water) during 20 seconds sharp. There is no need for agitation during etching. Rinsing is performed by a first bath of water during 20 seconds with agitation, then in a second bath of water for tranport between benches.

### v2

@@ IV.2 Four-wave-mixing tuning
The 4WM condition is found by sending a pump tone at frequency $\omega_{p}$ together with a signal at frequency $\omega$ into the SMPD input, followed by qubit readout. Sweeping $\omega_{p}$ and $\omega$, the 4WM condition is identified by an enhanced qubit excited state when the signal is present, which vanishes when it is switched off (see Fig. 7) ; the pump power is then tuned to reach unit cooperativity (see Appendix and Fig. 7).
@@ IV.3 Cyclic operation : Detector bandwidth and dark counts
We now turn to the operation of the SMPD in cyclic mode. The detection step duration is set to $T_{d}=15$ µs. When no photon is sent to the device, the qubit is most often in the ground state at the end of the sequence, implying that no reset is needed, and that the average cycle duration is very close to $T_{d}+T_{\mathrm{RO}}=15.8$ µs, yielding a duty cycle $\eta_{\mathrm{cycle}}=0.95$. From Eq. 3, we expect $\eta_{\mathrm{SMPD}}$ to reach $0.74$.
@@
The experimental $\eta_{\mathrm{SMPD}}$ is measured by counting the number of clicks detected over 1 second integration time, at the same time that a calibrated signal is sent to the SMPD input. The signal power is calibrated using the qubit as a detector (see [4] and Appendix B). The result is plotted as a function of input frequency $\omega$ in Fig. 2d. The maximum efficiency reaches $\eta_{\mathrm{SMPD}}=0.8$. This is even slightly higher than the predicted value, which implies in particular that the 4WM efficiency $\eta_{\mathrm{4WM}}$ is very close to $1$. Fitting $\eta_{\mathrm{SMPD}}(\omega)$ with a lorentzian, we obtain the detector bandwidth $\kappa_{d}$. Two curves are shown, obtained for two different values of $\kappa_{b}$. The maximum efficiency is unchanged, but the detector bandwidth is changed according to $\kappa_{b}$. This confirms that the SMPD bandwidth can be adjusted by tuning $\kappa_{b}$ via the Purcell filter frequency. We measure $\kappa_{d}/2\pi=315$ kHz for $\kappa_{b}/2\pi=170$ kHz, and $\kappa_{d}/2\pi=530$ kHz for $\kappa_{b}/2\pi=280$ kHz, thus confirming the approximate relation $\kappa_{d}\simeq 2\kappa_{b}$ (see Appendix A).
@@
Dark counts are measured by recording photon-counting traces in the absence of any input signal (see Tab. 3, where the detector bandwidth was $\kappa_{d}/2\pi=170$ kHz). To obtain the various contributions to the dark counts, the traces are systematically measured with the pump off (yielding $\alpha_{q}$), with the pump on and detuned by $20$ MHz (yielding $\alpha_{q}+\alpha_{p}=\alpha_{\mathrm{err}}$), and with the pump on and tuned to the 4WM conditions (yielding $\alpha=\alpha_{\mathrm{err}}+\alpha_{\mathrm{th}}$). At $10$ mK, we obtain $\alpha_{q}=8\mathrm{s}^{-1}$, $\alpha_{p}=2\mathrm{s}^{-1}$, and $\alpha_{\mathrm{th}}=21\mathrm{s}^{-1}$. Combining this with the measured efficiency, we obtain a power sensitivity $S=3\cdot 10^{-23}\mathrm{W}/\sqrt{\mathrm{Hz}}$, a factor of 3 better than the previous state of the art [4], which corresponds to an order of magnitude gain in measurement time.
@@
Time traces are similarly measured by varying the cryostat temperature $T$. The values of $\alpha_{q}$ and $\alpha_{\mathrm{th}}$ are shown as a function of $T$ in Fig. 3. The qubit contribution increases with temperature because of the higher excited-state probability at equilibrium; it is well fitted by $\alpha_{q,0}+K_{q}p_{\mathrm{th},q}(T)$, with $\alpha_{q,0}=7\,\mathrm{s}^{-1}$ and $K_{q}=2.2\cdot 10^{4}\,\mathrm{s}^{-1}$. The value of $K_{q}$ is slightly larger than the predicted $(T_{d}/T_{1})(1/(T_{d}+T_{\mathrm{RO}}+T_{\mathrm{reset}}))$ (see Section IIIB), which is $1.4\cdot 10^{4}\,\mathrm{s}^{-1}$, possibly because of a reduction of $T_{1}$ at higher temperature. The thermal contribution also increases with temperature; it is well fitted by $\alpha_{\mathrm{th},0}+K_{\mathrm{th}}\bar{n}_{\mathrm{th},b}$, with $\alpha_{\mathrm{th},0}=31\mathrm{s}^{-1}$, and $K_{\mathrm{th}}=2\cdot 10^{5}\,\mathrm{s}^{-1}$. This value is in quantitative agreement with the expected $\eta_{\mathrm{SMPD}}\kappa_{d}/4=2.1\cdot 10^{5}\mathrm{s}^{-1}$, thus providing an independent confirmation of the efficiency estimate reported earlier.
@@
The equilibrium qubit excitation (without active reset) $p_{\mathrm{th},q}$ at $10$ mK translates into an effective qubit temperature of $42$ mK. The thermal photon rate $\alpha_{\mathrm{th},0}$ measured at $10$ mK translates into a mean photon number per mode of $1.5\cdot 10^{-4}$ and thus to an effective temperature of $40$ mK. This is higher than the mean photon number expected from our line attenuation, which we estimate to be below $\sim 2\cdot 10^{-5}$. Our results therefore confirm the well-known difficulty of thermalizing microwave fields and modes to temperatures much lower than $\sim 40$ mK, possibly due to the difficulty of perfectly thermalizing microwave components. It is interesting to note that SMPD devices are ideally well suited to studying this phenomenon, which represents a major issue for superconducting quantum computers.
@@
Table 3: Photon-counting traces taken under different conditions. First line : the mixing-chamber temperature is 10 mK and the pump is off, yielding $\alpha_{q}$. Second line : the mixing-chamber temperature is 10 mK and the pump is switched on, but it is detuned by $20$ MHz from the 4WM condition, yielding $\alpha_{\mathrm{err}}=\alpha_{q}+\alpha_{p}$. Third line : the mixing-chamber temperature is 10 mK and the pump is on and in the 4WM condition, yielding $\alpha_{\mathrm{err}}+\alpha_{\mathrm{th}}$. Three following lines : the mixing-chamber temperature is increased, and the count rate rises accordingly, as detailed in fig. 3.
@@ IV.4 Dependence on detection bandwidth
Figure 4: The blue full circles are the measured total dark counts, $\alpha_{\mathrm{err}}+\alpha_{\mathrm{th}}$, as a function of SMPD bandwidth $\kappa_{d}$. The dashed line is a linear fit, with slope $1.5\cdot 10^{-5}$, and intercept $11.6\mathrm{s}^{-1}$. The red triangles show the measured efficiency at resonance, $\eta_{\mathrm{SMPD}}$.
We finally study the SMPD figures of merit as a function of detection bandwidth, at the lowest temperature of $10$ mK. We measure the maximum value of $\eta_{\mathrm{SMPD}}(\omega)$ and $\alpha$ as a function of $\kappa_{d}$ (see Fig. 4). We observe that $\eta_{\mathrm{SMPD}}$ increases with $\kappa_{d}$, although the dependence is very weak as soon as the SMPD bandwidth is above $250$ kHz. This is expected because the measurements are performed with a monochromatic tone.The reduction in efficiency at the lowest value of $\kappa_{d}$ may be attributed to the contribution of internal buffer losses, as well as possible transient effects that effectively shorten the detection time $T_{d}$.
The false-positive rate $\alpha$, however, scales linearly with $\kappa_{d}$, as expected from Eq. 5. The fitted slope is $1.6\cdot 10^{-5}$, in agreement with the expected $\eta_{\mathrm{SMPD}}\bar{n}_{\mathrm{th},b}/4$ yielding a thermal mode occupation $\bar{n}_{\mathrm{th},b}=0.8\cdot 10^{-4}$, slightly lower than in the measurements of Fig. 3 and corresponding to a mode temperature of $38$ mK. The intercept at $\kappa_{d}=0$ is $11.6\,\mathrm{s}^{-1}$, consistent with the error rate $\alpha_{\mathrm{err}}$ measured with pump detuned. Overall, the observed linear dependence of $\alpha_{\mathrm{th}}$ on $\kappa_{d}$ proves that these counts are indeed due to thermal photons impinging on the SMPD input line. Not much can be inferred about their spectrum, apart from the fact that it appears to be flat on the scale of the detection bandwidth ($\sim 1$ MHz).
For fig. 4, no data have been taken above $\kappa_{d}/2\pi=$ 1 MHz. Indeed, above this value, high pump amplitudes are required to keep $C=1$, and our microwave setup was unable to send high enough power.
@@ IV.5 Single-spin measurements
We confirm the high efficiency of the detector by measuring the microwave fluorescence from a single electronic spin. The setup closely follows refs [41] and [39], and is schematically depicted in Fig. 5. The electronic spin is an $\mathrm{Er}^{3+}$ ion in a $\mathrm{CaWO}_{4}$ matrix and is coupled to a superconducting resonator to enhance its radiative decay rate $\Gamma_{R}$. The spin is excited by a $\pi$ pulse with near-unit probability. It then relaxes radiatively by emitting a photon, which is either absorbed by the internal losses of the spin resonator, or emitted in the line, after which it may reach the SMPD detector or be absorbed by the losses between the two devices. The SMPD buffer $\omega_{b}$ is tuned to the spin and resonator frequency. The spin-to-click probability is therefore $\eta=\eta_{reso}\eta_{loss}\eta_{\mathrm{SMPD}}(\omega_{b})$, with $\eta_{reso}=0.6$ in this device (described in more detail elsewhere [39]).
The spin microwave fluorescence curve is shown in Fig. 5. This is well fitted by an exponential decay with time constant $\Gamma_{R}^{-1}=1.24$ ms on top of a constant background, due to the SMPD dark counts. From the integral below the fluorescence curve, we obtain $\eta=0.4$. This implies that the microwave power absorption between the spin device and the SMPD is $\eta_{loss}=0.85$, a very plausible value given that the line includes a superconducting cable, a circulator, and an infrared filter. This provides an independent confirmation of the high efficiency reached by this SMPD device and of its relevance for single-spin spectroscopy.
@@ V Conclusion
We have demonstrated a microwave photon counter with tunable frequency (in the $7$ GHz range) and bandwidth (between $\sim 100$ kHz and $\sim 1$ MHz). The device has a maximum efficiency of $0.8$, and its noise can be reduced by lowering the bandwidth. At an input bandwidth of $200$ kHz, we demonstrate a record power sensitivity of $3\cdot 10^{-23}\mathrm{W}/\sqrt{\mathrm{Hz}}$. The device is well suited to detecting emitters with a narrow line, and it will find applications in single-spin magnetic resonance [39, 30] as well as in haloscope-based dark-matter searches [12].
@@ Acknowledgements
We acknowledge technical support from P. Sénat, D. Duet, P.-F. Orfila, and S. Delprat, and we are grateful for fruitful discussions within the Quantronics group. We acknowledge support from the Agence Nationale de la Recherche (ANR) through the MIRESPIN (ANR-19-CE47-0011) project. We acknowledge support of the Région Ile-de-France through the DIM QUANTIP, from the AIDAS virtual joint laboratory, and from the France 2030 plan under Grant No. ANR-22-PETQ-0003. This project has received funding from the European Union Horizon 2020 research and innovation program under the OpenSuperQ100+ project, and from the European Research Council under Grant No. 101042315 (INGENIOUS). We acknowledge IARPA and Lincoln Labs for providing the Josephson traveling-wave parametric amplifier.
Figure 5: Microwave fluorescence of individual $\mathrm{Er}^{3+}$ electron spins in a $\mathrm{CaWO}_{4}$ host crystal matrix at $10$ mK, detected by the SMPD reported here operated with $\kappa_{d}/2\pi=500$ kHz (blue) and by a previous-generation SMPD operated with $\kappa_{d}/2\pi=400$ kHz (red). Top : scheme of the experiment. The $Er^{3+}$ spin is magnetically coupled to the microwave field $\vec{B}_{1}$ of an LC resonator and Zeeman tuned to resonance by a field $\vec{B}_{0}||$. The spin frequency is set to 7.7492 GHz in the latest version of the experiment, and 7.335 GHz for the ”old” version. The spin is excited by a $\pi$-pulse, and the photon it emits upon radiative relaxation is detected by the SMPD. Graph : blue full circles are the measured count rate following a $\pi$ pulse excitation of the spin, as a function of the time, with the SMPD presented herein. The solid blue line is an exponential fit, from which we extract the spin lifetime $\Gamma_{R}^{-1}=1.24$ ms and spin-to-click efficiency, $\eta=0.4$. Note that the noise floor $\alpha_{err}+\alpha_{th}=0.12\mathrm{ms}^{-1}$ is slightly higher than expected from Fig. 4, possibly due to a different wiring scheme in the cryostat. The red full circles and solid line are the equivalent data and fit for the same experiment performed with a previous-generation SMPD. One can observe a fourfold improvement of $\eta$ from the previous to the newest SMPD generation, yielding a measurement time faster by an order of magnitude.
[…]
@@
Figure 7: a) Four-wave mixing pattern : this color plot shows the measured qubit excited state as a function of pump frequency $\omega_{p}$ and signal frequency $\omega$ (from the input line). The qubit is excited only when the signal is close to $\omega_{b}$, and the pump verifies the 4WM condition. b) Fit from a two-coupled-cavity model. To fit the data, the value of the cooperativity $C$ is adjusted. Here, the value is $C=0.99$ (see appendix A). Other fit parameters are the buffer linewidth, $\nicefrac{{\kappa_{b}}}{{2\pi}}=120$ kHz, and the waste linewidth, $\nicefrac{{\kappa_{w}}}{{2\pi}}=1.75$ MHz. c) Same experiment as in a, but with the signal tone switched off. The higher excitation probabilities seen in subfigure a are no longer visible, demonstrating that we were indeed observing the actual 4WM pattern. Insets : corresponding pulse sequences. Purple pulses are sent through the pump/drive line, orange pulses through the signal/input line and green pulses through the waste/readout (RO) line. The pumping, signal, and readout pulses are respectively 15 µs, 10.5 µs and 800 ns long.
In this appendix we detail the tuning procedure for the pump parameters. The optimal pump frequency should satisfy the 4WM condition, $\omega_{\mathrm{4WM}}=\omega_{q}+\omega_{w}-\chi_{w}-\omega_{b}$. In the calibration procedure of the device, pump calibration comes after qubit characterization and readout calibration, which results are displayed in Fig. 6.
The pump amplitude is chosen by recording two-dimensional plots such as those shown in Fig. 7, where the qubit excited-state probability is recorded as a function of pump and signal frequency, $\omega_{p}$ and $\omega$, respectively, for various pump powers. Each plot is then fitted to the following formula [2] :
@@
where $\delta=\omega-\omega_{b}$ is the signal-buffer detuning, and $\delta_{p}=\omega_{p}-\omega_{\mathrm{4WM}}$ the pump detuning from the 4WM condition.
@@
Figure 8: Detection bandwidth $\kappa_{d}$ as a function of buffer bandwidth $\kappa_{b}$. Full blue circles are experimental data and the blue solid line is a linear fit. The fit returns a slope of 1.9, which is close to 2, the value that is expected from eq. 9
Fig. 8 shows that eq. 9 is verified experimentally.
@@
Measuring the efficiency of the SMPD begins with estimating the power at the buffer input. This is achieved by using the qubit itself as a detector. Ramsey fringes are shifted by $\Delta_{q}$ and dephased by $\Gamma_{q}$ according to the equation:
@@
where $\epsilon$ is the complex amplitude of the intracavity field. The efficiency is then simply computed by dividing the number of clicks per second by the number of photons per second at the device input.
@@
The process starts with a two-inch, double-side-polished sapphire wafer. This wafer is dipped in a piranha solution made with 2/3 sulfuric acid and 1/3 hydrogen peroxyde (by volume) and left for 20 minutes. Then, it is rinsed with water and dried with a stream of nitrogen. After that, the wafer is loaded into an RF reactive sputterer together with a tantalum target. The targeted thickness of the tantalum sputtering is between 80 and 90 nm. The sputtering process shall be performed while the wafer is heated up to 600°C to obtain alpha phase tantalum.
The wafer is then prepared for dicing. To this end, UV III is spin-coated over the wafer for 60 seconds at 3000 rpm with an acceleration of 2000 rpm/s, followed by a postbake at 120°C for 3 minutes.
@@
After dicing, the protective UV III layer is removed applying isopropanol (IPA) from a wash bottle the chosen chip for 15 seconds, then again for 5 seconds. For further cleaning, the chip is immersed in acetone under strong ultrasonic agitation for 5 minutes, then in another ultrasonic bath of IPA for 3 minutes. In the following, it is then dried with a stream of nitrogen and immersed in a piranha solution for 20 minutes. The chip is then rinsed with water, dehydrated with IPA, dried with a stream of nitrogen, and further dehydrated on a hotplate at 120°C for 5 minutes. Next, AZ1518 is spin-coated for 60 seconds at 3000 rpm with an acceleration of 2000 rpm/s. The postbake is performed on a hotplate for 2 minutes at 115°C. The circuit is then patterned using a UV LASER direct writer. A supplementary exposure is needed at the edge bead. The development is performed in pure MF319 for 60 seconds and stopped with water. The chip is dried with a stream of nitrogen and hard baked on a hotplate for 2 minutes at 115°C. Finally, the circuit is etched by dipping in Transene 111 (1/3 nitric acid, 1/3 hydrofluoric acid, and 1/3 water) for exactly 20 seconds. There is no need for agitation during etching. Rinsing is performed in a first bath of water for 20 seconds with agitation and then in a second bath of water for transport between benches.